OpenStudy (anonymous):

simplify the expression n^2-n Over n^2-5n

OpenStudy (anonymous):

lets factorize \(n^2-n\) first what is the common factor between \(n\) and \(n^2\) ?

OpenStudy (anonymous):

n?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok so the top half is n

OpenStudy (anonymous):

what will be the top after taking n as common?

OpenStudy (anonymous):

I do not know

OpenStudy (anonymous):

\[n^2-n=n\times n -n=n(n-1)\]